# A Non-Relativistic Limit for Heterotic Supergravity and its Gauge Lagrangian

### v3

@@ 3.2 Gauge transformations
Transformation rules under the non-Abelian gauge symmetry can be extracted from (1.9), (1.10) and (1.11). We will impose an extra expansion on the gauge parameter and the structure constants following [22],
@@
At this point, one can think that the rescaling of the structure constants might include divergences in the gauge Lagrangian. However, this reescaling is safe because higher order $\hat{f}_{ijk}$ contributions are $\propto\frac{1}{c}$. Then, the rescaling only introduces cubic contributions in gauge fields in the finite Lagrangian, which we will present in the next section.
After exploring the gauge transformation (1.9), (1.10) and (1.11) we find the following rules upon taking the limit $c\rightarrow\infty$,
@@
$\delta_{\lambda}b_{\mu\nu}$ $=$ $-(\partial_{[\mu}\lambda^{i})\alpha_{-i}\tau_{\nu]}^{-}\,,$ (3.11)
$\delta_{\lambda}h_{\mu\nu}$ $=$ $0\,,$ (3.12)
$\delta_{\lambda}h^{\mu\nu}$ $=$ $0\,,$ (3.13)
$\delta_{\lambda}\tau_{\mu}{}^{a}$ $=$ $0\,,$ (3.14)
$\delta_{\lambda}\tau^{\mu}{}_{a}$ $=$ $0\,.$ (3.15)
From the previous transformations, we learn that $a_{\mu}^{i}$ transforms as a gauge connection, $\alpha_{-}^{i}$ transforms as a vector and $b_{\mu\nu}$ transforms in a non-covariant way, emulating a Green-Schwarz mechanism. Interestingly, this mechanism cannot be eliminated through field redefinitions in the relativistic scenario. However, when one considers the NR limit, it is possible to define the field
$\bar{b}_{\mu\nu}=b_{\mu\nu}+a_{[\mu|i}\alpha_{-}^{i}\tau_{|\nu]-}$ (3.16)
which is gauge invariant, i.e. $\delta_{\lambda}\bar{b}_{\mu\nu}=0$. This is an interesting property of the heterotic supergravity under the NR limit which will help us to construct a suitable curvature for the $\bar{b}$-field.
@@ 3.3 Gauging the covariant derivative
The field $\alpha_{-}^{i}$ is already covariant, so we do not need to construct its curvature. If we consider $\bar{b}$ as the fundamental field of the theory, then we define its curvature as
$\bar{h}_{\mu\nu\rho}=3\partial_{[\mu}\bar{b}_{\nu\rho]}\,=h_{\mu\nu\rho}-3c^{(g)}_{\mu\nu\rho}.$ (3.19)
@@
In the next section, we will use these curvatures to present the finite NR gauge Lagrangian.
@@ 4 Finite Lagrangian
In this section we construct all the finite contributions to the heterotic supergravity Lagrangian in covariant form upon taking the NR limit. For the readers interested only in the final result, we report the final form of the gauge Lagrangian in the conclusions.
@@ 4.1 Ungauged Lagrangian
We would like to present the finite heterotic Lagrangian in terms of covariant quantities. We will choose objects which are covariant under infinitesimal diffeomorphisms. Due to the previous section, we have all the pieces to write the ungauged part of the finite Lagrangian in its covariant form. This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term and the H-term given by
$L$ $=$ $R^{(0)}+4\partial_{\mu}\varphi\partial^{\mu}\varphi-\frac{1}{12}H^{(0)}_{\mu\nu\rho}H^{(0)}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$ (4.1)
$-\frac{1}{2}H^{(0)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}H^{(2)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\eta^{cd}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c}\tau^{\sigma}{}_{b}\tau^{\gamma}{}_{d}h^{\rho\epsilon}\,,$
@@
$H^{(2)}_{\mu\nu\rho}$ $=$ $-6c^{2}\epsilon_{ab}\nabla_{[\mu}\tau_{\nu}{}^{a}\tau_{\rho]}{}^{b}\,,$ (4.3)
and the Ricci scalar is given by
$R^{(0)}$ $=$ $8\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau^{\mu}{}_{a}}h^{\nu\rho}+5\nabla_{[\mu}{\tau_{\nu]}{}^{c}}\eta^{ab}(\nabla_{\rho}{\tau^{\mu}{}_{a}}\tau^{\nu}{}_{c}\tau^{\rho}{}_{b}+\nabla_{\rho}{\tau^{\rho}{}_{a}}\tau^{\mu}{}_{b}\tau^{\nu}{}_{c})$ (4.4)
$-\frac{3}{8}\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}(\tau^{\nu}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\rho}-2\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\rho}+\tau^{\mu}{}_{a}\tau^{\rho}{}_{b}h^{\nu\sigma})$
$+\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}h^{\mu\rho}h^{\nu\sigma}+4c^{2}\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}h^{\mu\sigma}h^{\nu\rho}$
$+\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}\eta^{cd}(2\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}-\tau^{\mu}{}_{c}\tau^{\rho}{}_{d}h^{\nu\sigma}-\tau^{\nu}{}_{c}\tau^{\sigma}{}_{d}h^{\mu\rho})$
$+\frac{1}{c^{2}}\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}\eta^{cd}(\tau^{\nu}{}_{c}\tau^{\sigma}{}_{d}h^{\mu\rho}-2\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}+\tau^{\mu}{}_{c}\tau^{\rho}{}_{d}h^{\nu\sigma})$
$-2\nabla_{\mu}{\nabla_{\nu}{\tau_{\rho}{}^{a}}}\tau^{[\rho}{}_{a}h^{\nu]\mu}+f(Q,S)\,,$
where $f(Q,S)$ is a covariant contribution since both $Q_{\mu\nu}{}^{a}$ and $S_{\mu}{}^{\nu}{}_{a}$ are covariant,
$f(Q,S)$ $=$ $8Q_{\mu\nu}{}^{a}(S_{\rho}{}^{[\mu}{}_{a}h^{\rho]\nu}+\nabla_{\rho}{\tau^{[\rho}{}_{a}}h^{\mu]\nu})-8\nabla_{[\nu}{\tau_{\rho]}{}^{a}}S_{\mu}{}^{\nu}{}_{a}h^{\mu\rho}-\frac{1}{2}Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\tau^{\mu}{}_{a}\tau^{\rho}{}_{b}h^{\nu\sigma}$ (4.5)
$+Q_{\mu\nu}{}^{a}\eta^{bc}(-2\nabla_{\rho}{\tau^{\mu}{}_{b}}\tau^{\nu}{}_{c}\tau^{\rho}{}_{a}+\nabla_{\rho}{\tau^{\mu}{}_{b}}\tau^{\nu}{}_{a}\tau^{\rho}{}_{c}+\nabla_{\rho}{\tau^{\rho}{}_{b}}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c})$
$+Q_{\mu\nu}{}^{a}\eta^{bc}(2S_{\rho}{}^{\mu}{}_{b}\tau^{\nu}{}_{c}\tau^{\rho}{}_{a}-S_{\rho}{}^{\mu}{}_{b}\tau^{\nu}{}_{a}\tau^{\rho}{}_{c}-S_{\rho}{}^{\rho}{}_{b}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c})$
$+\nabla_{\rho}{\tau_{\sigma}{}^{b}}Q_{\mu\nu}{}^{a}(-\frac{3}{2}\tau^{\rho}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\nu}+2\tau^{\mu}{}_{a}\tau^{[\rho}{}_{b}h^{\sigma]\nu}+2\tau^{\rho}{}_{b}\tau^{\sigma}{}_{a}h^{\mu\nu}+4\tau^{\mu}{}_{b}\tau^{[\rho}{}_{a}h^{\sigma]\nu})$
$+\nabla_{\nu}{\tau_{\rho}{}^{c}}\eta^{ab}(-\frac{5}{2}S_{\mu}{}^{\nu}{}_{a}\tau^{\mu}{}_{b}\tau^{\rho}{}_{c}+5S_{\mu}{}^{\mu}{}_{a}\tau^{[\nu}{}_{c}\tau^{\rho]}{}_{b})$
$+\frac{5}{2}\nabla_{\rho}{\tau_{\nu}{}^{c}}\eta^{ab}S_{\mu}{}^{\nu}{}_{a}\tau^{\mu}{}_{b}\tau^{\rho}{}_{c}+2Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\eta_{ab}h^{\mu[\nu}h^{\rho]\sigma}+2Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\eta_{ab}\eta^{cd}\tau^{\mu}{}_{c}\tau^{[\rho}{}_{d}h^{\nu]\sigma}$
$+3\nabla_{\rho}{\tau_{\sigma}{}^{b}}Q_{\mu\nu}{}^{a}\eta_{ab}\eta^{cd}\tau^{\mu}{}_{c}\tau^{[\sigma}{}_{d}h^{\rho]\nu})+\frac{4}{c^{2}}\nabla_{[\rho}{\tau_{\sigma]}{}^{b}}Q_{\mu\nu}{}^{a}(\tau^{\rho}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\nu}+\tau^{\mu}{}_{b}\tau^{\sigma}{}_{a}h^{\nu\rho})$
$+4\nabla_{\mu}{Q_{\nu\rho}{}^{a}}\tau^{[\mu}{}_{a}h^{\nu]\rho}\,.$
In $R^{(0)}$ we can explicitly see that terms proportional to $c^{2}$ and $\frac{1}{c^{2}}$ naturally arises when we write this object using covariant derivatives. When we replace the covariant derivatives by their definitions, all these divergent and convergent quantities cancel, giving a finite contribution in terms of partial derivatives.

### v5

@@ 3.3 Gauge transformations
The transformation rules under the non-Abelian gauge symmetry can be extracted from (1.13) and (1.14). We will impose an extra expansion on the gauge parameter and the structure constants following [22],
@@
At this point, one can think that the rescaling of the structure constants might include divergences in the gauge Lagrangian. However, this rescaling is safe because higher order $\hat{f}_{ijk}$ contributions are $\propto\frac{1}{c}$. Then, the rescaling only introduces cubic contributions in gauge fields in the finite Lagrangian, which we will present in the next section.
After exploring the gauge transformations (1.13) and (1.14) we find the following rules when taking the limit $c\rightarrow\infty$,
@@
$\delta_{\lambda}b_{\mu\nu}$ $=$ $-(\partial_{[\mu}\lambda^{i})\alpha_{-i}\tau_{\nu]}^{-}\,,\,$, (3.16)
while the remaining fields do not transform under gauge transformations. From the previous transformations, we learn that $a_{\mu}^{i}$ transforms as a gauge connection, $\alpha_{-}^{i}$ transforms as a vector and $b_{\mu\nu}$ transforms in a non-covariant way, emulating a Green-Schwarz mechanism. While in the relativistic case this transformation is non-ambiguous and cannot be removed, when one considers the NR expansions presented in this work (based on [22]), it is possible to define the field
$\tilde{b}_{\mu\nu}=b_{\mu\nu}+a_{[\mu|i}\alpha_{-}^{i}\tau_{|\nu]-}$ (3.17)
which is gauge invariant, i.e. $\delta_{\lambda}\tilde{b}_{\mu\nu}=0$. This is an interesting property of the heterotic supergravity under the NR limit, which will help us to construct a suitable curvature for the $\tilde{b}-$field and, also, is potentially telling us that the cancellation of anomalies could be trivial under this limit. If this is the case, one could construct different formulations of heterotic supergravity using arbitrary gauge groups, beyond $SO(32)$ and $E8\times E8$.
The NR NS-NS supergravity is invariant under boost transformations given by [13],
$\delta_{\lambda}h_{\mu\nu}$ $=$ $-2\lambda_{aa^{\prime}}e_{(\mu}{}^{a^{\prime}}\tau_{\nu)}{}^{a}\,$, (3.18)
$\delta_{\lambda}\tau^{\mu\nu}$ $=$ $2\lambda_{a}{}^{a^{\prime}}e^{(\mu}{}_{a^{\prime}}\tau^{\nu)a}\,$, (3.19)
$\delta_{\lambda}b_{\mu\nu}$ $=$ $-2\epsilon_{ab}\lambda^{a}{}_{a^{\prime}}\tau_{[\mu}{}^{b}e_{\nu]}{}^{a^{\prime}}\,$. (3.20)
Since the gauge fields do not transform under this symmetry, the field redefinition (3.17) do not deform the boost transformation of the new $\tilde{b}_{\mu\nu}$ field.
@@ 3.4 Gauging the covariant derivative
The field $\alpha_{-}^{i}$ is already covariant, so we do not need to construct its curvature. If we consider $\tilde{b}$ as the fundamental field of the theory, then we define its curvature as
$\bar{h}_{\mu\nu\rho}=3\partial_{[\mu}\tilde{b}_{\nu\rho]}\,=h_{\mu\nu\rho}-3c^{(g)}_{\mu\nu\rho}$. (3.23)
@@
The introduction of the $\tilde{b}_{\mu\nu}$ field is a crucial result of this paper. It shows that the gauge Green-Schwarz mechanism is trivial in the NR-limit of the heterotic supergravity. Furthermore, the Chern-Simons terms in form language are given by the exterior derivative of a two form, simplifying the Bianchi identity when we compare to the relativistic case. In section (4), we will use these curvatures to present the finite NR gauge Lagrangian.
@@ 4 Finite Lagrangian
In this section, we construct all the finite contributions to the heterotic supergravity Lagrangian in covariant form upon taking the NR limit. For readers interested only in the final result, we report the final form of the gauge Lagrangian at the end of this section.
@@ 4.1 Ungauged Lagrangian
This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term, and the H-term, given by
$L$ $=$ $ee^{-2\varphi}(R^{(0)}+4\partial_{\mu}\varphi\partial^{\mu}\varphi-\frac{1}{12}H^{(0)}_{\mu\nu\rho}H^{(0)}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$ (4.1)
$-\frac{1}{2}H^{(0)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}H^{(2)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\eta^{cd}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c}\tau^{\sigma}{}_{b}\tau^{\gamma}{}_{d}h^{\rho\epsilon})\,$,
@@
$H^{(2)}_{\mu\nu\rho}$ $=$ $-6c^{2}\epsilon_{ab}\nabla_{[\mu}\tau_{\nu}{}^{a}\tau_{\rho]}{}^{b}\,$. (4.3)
Collecting all the contributions coming from the H-terms, we find the following finite Lagrangian
$L=ee^{-2\varphi}(R^{(0)}+4\partial_{\mu}\varphi\partial_{\nu}\varphi h^{\mu\nu}+T_{H^{2}})$ (4.4)
where
$T_{H^{2}}$ $=$ $2\eta_{ab}\eta^{cd}\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{[\rho}{\tau_{\sigma]}{}^{b}}\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}+2\epsilon_{ab}\epsilon_{cd}\eta^{ae}\eta^{cf}\nabla_{[\mu}{\tau_{\nu]}{}^{b}}\nabla_{[\rho}{\tau_{\sigma]}{}^{d}}\tau^{\nu}{}_{f}\tau^{\sigma}{}_{e}h^{\mu\rho}$ (4.5)
$-\frac{1}{12}h_{\mu\nu\rho}h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\epsilon_{ab}\eta^{ac}\nabla_{[\mu}{\tau_{\nu]}{}^{b}}h_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\rho}h^{\nu\gamma}$
The Ricci scalar can be easily computed considering the non-relativistic expansions (1.1)-(1.2). The generic form of this object is as follows,
$R=c^{2}R^{(2)}+R^{(0)}+\frac{1}{c^{2}}R^{-2}+\frac{1}{c^{4}}R^{-4}\,.$ (4.6)
Every contribution in the previous expansion transforms as a scalar, as the reader can easily check (we attach the explicit demostration for $R^{(0)}$ in the appendix C). The Lagrangian (4.4), together with (4.5) and $R^{(0)}$, constitutes the ungauged bosonic part of the heterotic supergravity.